OpenStudy (anonymous):

Factor the trinomial, or state that the trinomial is prime 7x^2+4x+3

OpenStudy (anonymous):

(x+1)(7x+2) or its prime... I think? Damn thats a tricky one.

OpenStudy (anonymous):

If I'm wrong its the +1 and the +2

jimthompson5910 (jim_thompson5910):

7x^2+4x+3 is prime Why? Multiply 7 and 3 to get 21. Now which two numbers multiply to 21 AND add to 4 (the middle coefficient)? There are no two rational numbers that do this, so the polynomial is prime.

OpenStudy (anonymous):

Wohoo I was right lol I felt kinda stupid.

OpenStudy (anonymous):

Or calculate the discriminant b^2-4ac and note that it is negative so no real solutions..

OpenStudy (anonymous):

If you want it factored, it does factor just not over the reals. It factors using the quadratic formula: \[\frac{-4 \pm \sqrt{16-4(7)(3)}}{2(7)}=\frac{-4 \pm \sqrt{-68}}{14}=\frac{-2 \pm \sqrt{17}i}{7}\] So it factors to: \[(x-(\frac{-2+\sqrt{17}i}{7}))(x-(\frac{-2-\sqrt{17}i}{7}))\]

OpenStudy (anonymous):

can you help with my other Q'S?

OpenStudy (anonymous):

Sure :P

OpenStudy (anonymous):

Factor and simplify the algebraic expression. \[(x+9)^1/4+(x+9_)^3/4\]